A beach ball is filled with 10.0 L of air in a hotel room where the temperature is 65 °F. What is the volume of the air in the beach ball out on the beach where the temperature is 38 °C?


1
Expert's answer
2021-03-29T05:19:03-0400

V1/T1 = V2/T2

65°F = 291.48 K

38°C = 308.15 K

10/291.48 = V2/308.15

V2 = (10/291.48) x 308.15 = 10.57L